# Provable Low-Frequency Bias of In-Context Learning of Representations

### v1

@@ 6.3 Why Are Peripheral Nodes Compressed?
In Park et al. (2024), the authors keenly observed that the empirical PCA results (say, for the grid graph as in Figure 2(c)), despite roughly showing the underlying grid structure, appears slightly compressed near the periphery, compared to the actual grid formed by the eigenvectors of the original graph (as in Figure 2(a), or Figure 7 in Park et al. (2024)). The authors attributed this distortion to uneven visitation frequencies in the random walk:
@@
While there is indeed a difference in visitation frequency, we argue that it is not the most fundamental explanation. The true cause lies in the context-wise process. As shown in Theorem 1, the transformation applied by the attention map to the latent representations is modulated by the stationary distribution ${\bm{\pi}}$. As a result, the actual graph the model is aware of is the reweighted graph ${\bm{W}}$ instead of the original one $\widetilde{\bm{W}}$. The top eigenvectors are twisted a little bit according to node degrees since it is reweighted by ${\bm{\pi}}$.
Figure 2: Comparison between empirical observations and theoretical predictions. (a): The 2nd and 3rd eigenvectors of $\widetilde{\bm{W}}$, illustrating the idealized grid structure hypothesized in Park et al. (2024); (b): The 2nd and 3rd eigenvectors of ${\bm{M}}$, as predicted by Theorem 2; (c): The first two principal components of the actual hidden representations from a pre-trained Transformer (Llama-3.1-8B), collected at context positions 2360-2560. In all subfigures, each point represents the latent representation of a word in the vocabulary. The $x$- and $y$-axes represent the values of the 2nd and 3rd eigenvectors (or the 1st and 2nd principal components), respectively. A small rotation was applied to panels (a) and (b) to better visually align them with (c). This is valid since the 2nd and 3rd eigenvalues of ${\bm{W}}$ are equal, and their eigenspace is invariant under rotation.
@@ 6.4 Why Energy Decreases but Doesn’t Vanish?
In Park et al. (2024), the authors hypothesized that the structure of the representation is a consequence of energy minimization. While this observation aligns with empirical trends, we argue that energy decay is not the fundamental cause, since 1) it doesn’t explain why the model follows the principle of energy decaying, and 2) the energy does not actually converges to 00, despite the 00-energy solutions are actually easy to find. Instead, both the energy decay and the structured representations are consequences of the double convergence.
@@
As shown in Theorem 2, the projections of the latent representations onto the low eigenspaces converges to ${\bm{0}}$ [^3] In principle, Theorem 2 is a relative result. However, with a similar proof one can show that the numerator also actually converge to 00 as long as the corresponding eigenvalues are significantly smaller than $1$., and thus $\left\|\Pi_{i}({\bm{z}}_{x}-{\bm{z}}_{y})\right\|$ converges to 00 for large $i$. The energy decay is thus a consequence of the representation leaving corresponding eigenspace.
@@
We validate this explanation empirically in Figure 4. While the overall energy decreases across layers, the energy in low-frequency directions (e.g., Component 1 and 2) increases, confirming our theoretical prediction: energy decay arises from the attenuation of high-frequency components, whereas the persistence of low-frequency components prevents the total energy from converging to 00.
@@ 6.5 Predicted Robustness Against Noise
Figure 3: The Normalized energy evolution across layers. Each “Component” curve shows the energy along a direction defined by the $k$-th eigenvector of ${\bm{M}}$. We only show the first $5$ components as an illustration. The “Overall” curve represents the average energy across all directions. To eliminate scale effects, the matrix ${\bm{Z}}^{(\ell)}$ is normalized to have unit Frobenius norm at each layer.
@@
To test this prediction, we conduct an additional experiment under a noisy data-generating process. Specifically, during the random walk over the graph $\mathcal{G}$, we inject noise by allowing the sequence to transition to a uniformly random token in $[c]$ with $1\%$ probability at each step, rather than to a graph neighbor. This corruption can be viewed as temporarily replacing the original graph with a complete graph, which introduces purely high-frequency components into the sequence.
In Figure 4, we measure the number of non-neighbor transitions (i.e. token pairs that do not correspond to valid edges in $\mathcal{G}$) within a sliding window of the last 500 tokens. We compare this quantity in both the input sequence and the output predicted by the ICL model. While the input maintains a constant error rate due to the injected corruption, the ICL output gradually eliminates these errors. Once the context becomes sufficiently long, the model consistently produces transitions that respect the original graph structure, effectively achieving $100\%$ accuracy despite the noisy input.
@@ 7 Summary
In this section, we prove Theorem 1. The proof start by identifying a high-probability event in the random walk sequence that ensures it is “regular” enough.
@@ A.1 Events in a Random Walk Sequence
Notice that in theorem 3, we view the spectral property of the graph as constant and absorb it into $C$. Below is a direct corollary of Theorem 3.
@@
in other words, with probability at least $0.999$, we have
@@
Notice that we assumed $n>10c$, therefore the following statement is also a direct corollary of Theorem 3.
@@
The following statement holds with probability at least $0.999$: for any $x\in[c]$, $F_{x,n}\geq F_{x,\lceil n/2\rceil}+1$.
A.2 Proof of Theorem 1
Now we prove Theorem 1. We start by a lemma that is easy to verify.
@@ A.2 Proof of Theorem 1
Subtracting eq. 27 into the argument above, we obtain
@@
Define $a_{k,y}=F_{y,k}/k$, $b_{k,y}=R/k$, $r_{y}=\pi_{y}$ and $s_{y}=\sum_{y^{\prime}\in f(x)}\pi_{y^{\prime}}$. From Corollary 4 we have there exists a constant number $C>0$ that only depends on $\mathcal{G}$, and an event whose probability is at least $0.999$, such that for all $k\in[n]$ and $y\in[c]$ we have $\max\left\{|a_{y,k}-r_{y}|,|b_{k,y}-s_{y}|\right\}\leq\frac{C\log n}{\sqrt{k}}$ (notice that this event is only related to the random walk, and does not depend on the specific values of $\mathcal{V}$, ${\bm{A}}$, etc.).
Let $\rho=\min_{y\in[c]}\pi_{y}\in(0,1)$. Let $C^{\prime}=2C/\rho>C$ that also only depends on $\mathcal{G}$. If $k\geq\frac{4C^{2}(\log n)^{2}}{\rho^{2}}$, then $\frac{C\log n}{\sqrt{k}}\leq\frac{\rho}{2}\leq\frac{s_{y}}{2}$, thus from Lemma 6 we have $\left|\frac{a_{y,k}}{r_{y}}-\frac{b_{y,k}}{s_{y}}\right|\leq\frac{C\log n}{\sqrt{k}}\leq\frac{C^{\prime}\log n}{\sqrt{k}}$. On the other hand, if $k<\frac{4C^{2}(\log n)^{2}}{\rho^{2}}$, we have $\frac{C^{\prime}\log n}{\sqrt{k}}\geq 1\geq\left|\frac{a_{y,k}}{r_{y}}-\frac{b_{y,k}}{s_{y}}\right|$. Therefore, we conclude that, in an event at happens with probability at least $0.999$, for all $k\in[n]$ and $y\in[c]$ we have $\left|\frac{F_{y,k}}{R}-\frac{\pi_{y}}{\sum_{y^{\prime}\in f(x)}\pi_{y^{\prime}}}\right|\leq\frac{C^{\prime}\log n}{\sqrt{k}}$.
Combining the above arguments, we conclude that with probability at least $0.999$, it holds that for all $k$,
@@
We first prove that under the specific conditions given in Section 5, how does the latent representations evolves.
@@
There exists a scalar number $C>0$ that possibly depends on the graph $\mathcal{G}$, and an event with probability at least $0.999$, such that the following statement holds. For any layer $\ell$, if ${\mathcal{V}}=\left\{{\bm{v}}_{k}\right\}_{k=1}^{n}$ is a good sequence converging to ${\mathcal{Z}}=\left\{{\bm{z}}_{x}\right\}_{x\in[c]}$ with parameter $\gamma$, and let ${\bm{A}}^{(\ell)}$ be defined as in eq. 7, then ${\mathcal{U}}={\bm{A}}^{(\ell)}{\mathcal{V}}$ is a good sequence converging to ${\bm{Z}}^{\prime}=\left\{{\bm{z}}^{\prime}_{x}\right\}_{x\in[c]}$, where
@@
Therefore, from Lemma 7 and Theorem 1, we have $\widehat{\mathcal{U}}$ converges to $\widehat{\mathcal{Z}}^{\prime}$ with parameter $\kappa$.
@@ B.1 Evolution of the Latent Representation
Notice that, since ${\bm{E}}$ is a rank-$1$ matrix, its image space is $\mathop{\mathrm{span}}\widetilde{\bm{1}}$: for any vector ${\bm{x}}\in\mathbb{R}^{c}$,
@@
From Lemma 9, we have
@@
Under the same condition as in Corollary 10, let $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a great mapping w.r.t. $\left(\gamma_{1},\gamma_{2},\left\{{\bm{Z}}^{\prime}\right\},{\bm{X}}\right)$ and $\left(\gamma_{1}^{\prime},\gamma_{2}^{\prime},\left\{{\bm{Z}}^{\prime}\right\},{\bm{Y}}\right)$. Then, we have
@@
Only need to repeat the proof of Corollary 10 and use the definition of great mappings. Notice that
@@
B.2 Proof of Theorem 2
Let $E_{1}$ be an event with probability at least $0.999$ defined in Lemma 8 holds. Let $E_{2}=\{\forall x\in[c],n_{[x]}>n/2\}$, Corollary 5 shows that $E_{2}$ also holds with probability at least $0.999$. In the following, we condition on the event $E_{1}\cap E_{2}$, which holds with probability at least $0.99$.
@@ B.2 Proof of Theorem 2
for any $\ell\in[L-1]$, and ${\bm{z}}_{x}^{\prime(\ell)}=\sigma^{(\ell)}\left({\bm{z}}_{k}^{(\ell)}\right)$. From the definition, it is obvious that $\left\{{\bm{v}}_{k}^{(1)}\right\}_{k=1}^{n}$ converges to $\mathcal{Z}^{(1)}$ with parameter 00.
  • We first fix an $L\in\mathbb{N}$ and analyze the context-wise convergence. First consider an arbitrary $x\in[c]$. Using Lemma 8 with an induction, it is not hard to prove that for each $\ell\in L$ there exists a $\gamma^{(\ell)}=\mathop{\mathrm{poly}}\log n$ (since we are going to take limit w.r.t. $n$, we view all other values independent of $n$ as constants; notice that $\ell$ is a fixed index here), such that $\left\{{\bm{v}}_{k}^{(\ell)}\right\}_{k=1}^{n}$ is a good sequence converging to $\left\{{\bm{z}}_{y}^{\prime(\ell)}\right\}_{y\in[c]}$ with parameter $\gamma^{(\ell)}$. Therefore, we have
@@
  • Next, we consider the layer-wise evolution. Corollary 11 shows that for any $\ell\in[L]$,
@@
    Corollary 11 also confirms that $\left\|{\bm{Z}}^{\prime(L)}{\bm{D}}^{1/2}{\bm{X}}\right\|>0$. Using an easy induction, we have
@@ Appendix C The Role of FFN: What Mappings are Great Mappings
The proof of Lemma 12 is obvious.
@@
The proof of Lemma 13 is obvious.
@@ Appendix D A Generalized Framework That is Independent of Data Generating
In the main paper, Theorem 1 conditions on the specific data generating process used in Park et al. (2024). This is because we need to match the distribution of the attention connection to each token with their stationary distribution in the data. In this section, we show that, with a slightly stronger assumption on the attention map, a general result of the context-wise convergence can be derived.
@@
Notice that Definition 6 is basically Definition 3 but limits the function $f$ maps each node to only one node, instead of a set of nodes as in Definition 3 (and in this case eq. 5 automatically holds). Although this condition seems stronger, we note that following the same idea used in the main paper, that we can compose multiple attention maps into one, this still represents a large family of allowed attention maps.
Next, we prove a similar result as Theorem 1 under Definition 6 that is independent of input distribution.
@@
Using, Lemma 7 which is also independent of DGP, we can prove the generalized results for a large family of attention maps by combining multiple attention maps that satisfies Definition 6.

### v2

@@ 6.3 Why Are Peripheral Nodes Compressed?
In Park et al. (2024), the authors keenly observed that the empirical PCA results (say, for the grid graph as in Figure˜2(c)), despite roughly showing the underlying grid structure, appears slightly compressed near the periphery, compared to the actual grid formed by the eigenvectors of the original graph (as in Figure˜2(a), or Figure 7 in Park et al. (2024)). The authors attributed this distortion to uneven visitation frequencies in the random walk:
@@
While there is indeed a difference in visitation frequency, we argue that it is not the most fundamental explanation. The true cause lies in the context-wise process. As shown in Theorem˜1, the transformation applied by the attention map to the latent representations is modulated by the stationary distribution ${\bm{\pi}}$. As a result, the actual graph the model is aware of is the reweighted graph ${\bm{W}}$ instead of the original one $\widetilde{\bm{W}}$. The top eigenvectors are twisted a little bit according to node degrees since it is reweighted by ${\bm{\pi}}$.
Figure 2: Comparison between empirical observations and theoretical predictions. (a): The 2nd and 3rd eigenvectors of $\widetilde{\bm{W}}$, illustrating the idealized grid structure hypothesized in Park et al. (2024); (b): The 2nd and 3rd eigenvectors of ${\bm{M}}$, as predicted by Theorem˜2; (c): The first two principal components of the actual hidden representations from a pre-trained language model (llama-3.1-8B), collected at context positions 2360-2560. In all subfigures, each point represents the latent representation of a word in the vocabulary. The $x$- and $y$-axes represent the values of the 2nd and 3rd eigenvectors (or the 1st and 2nd principal components), respectively. A small rotation was applied to panels (a) and (b) to better visually align them with (c). This is valid since the 2nd and 3rd eigenvalues of ${\bm{W}}$ are equal, and their eigenspace is invariant under rotation.
@@ 6.4 Why Energy Decreases but Doesn’t Vanish?
In Park et al. (2024), the authors hypothesized that the structure of the representation is a consequence of energy minimization. While this observation aligns with empirical trends, we argue that energy decay is not the fundamental cause, since 1) it doesn’t explain why the model follows the principle of energy decaying, and 2) the energy does not actually converges to 0, despite the 0-energy solutions are actually easy to find. Instead, both the energy decay and the structured representations are consequences of the double convergence.
@@
As shown in Theorem˜2, the projections of the latent representations onto the low eigenspaces converges to ${\bm{0}}$ [^4] In principle, Theorem 2 is a relative result. However, with a similar proof one can show that the numerator also actually converge to 0 as long as the corresponding eigenvalues are significantly smaller than 111., and thus $\left\|\Pi_{i}({\bm{z}}_{x}-{\bm{z}}_{y})\right\|$ converges to 0 for large $i$. The energy decay is thus a consequence of the representation leaving corresponding eigenspace.
@@
We validate this explanation empirically in Figure˜4. While the overall energy decreases across layers, the energy in low-frequency directions (e.g., Component 1 and 2) increases, confirming our theoretical prediction: energy decay arises from the attenuation of high-frequency components, whereas the persistence of low-frequency components prevents the total energy from converging to 0.
@@ 6.5 Predicted Robustness Against Noise
Figure 3: The Normalized energy evolution across layers. Each “Component” curve shows the energy along a direction defined by the $k$-th eigenvector of ${\bm{M}}$. We only show the first 555 components as an illustration. The “Overall” curve represents the average energy across all directions. To eliminate scale effects, the matrix ${\bm{Z}}^{(\ell)}$ is normalized to have unit Frobenius norm at each layer.
@@
[…]
@@ 7 Summary
In this section, we prove Theorem˜1. The proof start by identifying a high-probability event in the random walk sequence that ensures it is “regular” enough.
@@ A.1 Events in a Random Walk Sequence
Notice that in theorem˜3, we view the spectral property of the graph as constant and absorb it into $C$. Below is a direct corollary of Theorem˜3.
@@
in other words, with probability at least 0.9990.9990.999, we have
@@
Notice that we assumed $n>10c$, therefore the following statement is also a direct corollary of Theorem˜3.
@@
The following statement holds with probability at least 0.9990.9990.999: for any $x\in[c]$, $F_{x,n}\geq F_{x,\lceil n/2\rceil}+1$.
A.2 Proof of Theorem˜1
Now we prove Theorem˜1. We start by a lemma that is easy to verify.
@@ A.2 Proof of Theorem˜1
Subtracting eq.˜27 into the argument above, we obtain
@@
Define $a_{k,y}=F_{y,k}/k$, $b_{k,y}=R/k$, $r_{y}=\pi_{y}$ and $s_{y}=\sum_{y^{\prime}\in f(x)}\pi_{y^{\prime}}$. From Corollary˜4 we have there exists a constant number $C>0$ that only depends on $\mathcal{G}$, and an event whose probability is at least 0.9990.9990.999, such that for all $k\in[n]$ and $y\in[c]$ we have $\max\left\{|a_{y,k}-r_{y}|,|b_{k,y}-s_{y}|\right\}\leq\frac{C\log n}{\sqrt{k}}$ (notice that this event is only related to the random walk, and does not depend on the specific values of $\mathcal{V}$, ${\bm{A}}$, etc.).
Let $\rho=\min_{y\in[c]}\pi_{y}\in(0,1)$. Let $C^{\prime}=2C/\rho>C$ that also only depends on $\mathcal{G}$. If $k\geq\frac{4C^{2}(\log n)^{2}}{\rho^{2}}$, then $\frac{C\log n}{\sqrt{k}}\leq\frac{\rho}{2}\leq\frac{s_{y}}{2}$, thus from Lemma˜6 we have $\left|\frac{a_{y,k}}{r_{y}}-\frac{b_{y,k}}{s_{y}}\right|\leq\frac{C\log n}{\sqrt{k}}\leq\frac{C^{\prime}\log n}{\sqrt{k}}$. On the other hand, if $k<\frac{4C^{2}(\log n)^{2}}{\rho^{2}}$, we have $\frac{C^{\prime}\log n}{\sqrt{k}}\geq 1\geq\left|\frac{a_{y,k}}{r_{y}}-\frac{b_{y,k}}{s_{y}}\right|$. Therefore, we conclude that, in an event at happens with probability at least 0.9990.9990.999, for all $k\in[n]$ and $y\in[c]$ we have $\left|\frac{F_{y,k}}{R}-\frac{\pi_{y}}{\sum_{y^{\prime}\in f(x)}\pi_{y^{\prime}}}\right|\leq\frac{C^{\prime}\log n}{\sqrt{k}}$.
Combining the above arguments, we conclude that with probability at least 0.9990.9990.999, it holds that for all $k$,
@@
We first prove that under the specific conditions given in Section˜5, how does the latent representations evolves.
@@
There exists a scalar number $C>0$ that possibly depends on the graph $\mathcal{G}$, and an event with probability at least 0.9990.9990.999, such that the following statement holds. For any layer $\ell$, if ${\mathcal{V}}=\left\{{\bm{v}}_{k}\right\}_{k=1}^{n}$ is a good sequence converging to ${\mathcal{Z}}=\left\{{\bm{z}}_{x}\right\}_{x\in[c]}$ with parameter $\gamma$, and let ${\bm{A}}^{(\ell)}$ be defined as in eq.˜7, then ${\mathcal{U}}={\bm{A}}^{(\ell)}{\mathcal{V}}$ is a good sequence converging to ${\bm{Z}}^{\prime}=\left\{{\bm{z}}^{\prime}_{x}\right\}_{x\in[c]}$, where
@@
Therefore, from Lemma˜7 and Theorem˜1, we have $\widehat{\mathcal{U}}$ converges to $\widehat{\mathcal{Z}}^{\prime}$ with parameter $\kappa$.
@@ B.1 Evolution of the Latent Representation
Notice that, since ${\bm{E}}$ is a rank-111 matrix, its image space is $\mathop{\mathrm{span}}\widetilde{\bm{1}}$: for any vector ${\bm{x}}\in\mathbb{R}^{c}$,
@@
From Lemma˜9, we have
@@
Under the same condition as in Corollary˜10, let $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a great mapping w.r.t. $\left(\gamma_{1},\gamma_{2},\left\{{\bm{Z}}^{\prime}\right\},{\bm{X}}\right)$ and $\left(\gamma_{1}^{\prime},\gamma_{2}^{\prime},\left\{{\bm{Z}}^{\prime}\right\},{\bm{Y}}\right)$. Then, we have
@@
Only need to repeat the proof of Corollary˜10 and use the definition of great mappings. Notice that
@@
B.2 Proof of Theorem˜2
Let $E_{1}$ be an event with probability at least 0.9990.9990.999 defined in Lemma˜8 holds. Let $E_{2}=\{\forall x\in[c],n_{[x]}>n/2\}$, Corollary˜5 shows that $E_{2}$ also holds with probability at least 0.9990.9990.999. In the following, we condition on the event $E_{1}\cap E_{2}$, which holds with probability at least 0.990.990.99.
@@ B.2 Proof of Theorem˜2
for any $\ell\in[L-1]$, and ${\bm{z}}_{x}^{\prime(\ell)}=\sigma^{(\ell)}\left({\bm{z}}_{k}^{(\ell)}\right)$. From the definition, it is obvious that $\left\{{\bm{v}}_{k}^{(1)}\right\}_{k=1}^{n}$ converges to $\mathcal{Z}^{(1)}$ with parameter 0.
  • We first fix an $L\in\mathbb{N}$ and analyze the context-wise convergence. First consider an arbitrary $x\in[c]$. Using Lemma˜8 with an induction, it is not hard to prove that for each $\ell\in L$ there exists a $\gamma^{(\ell)}=\mathop{\mathrm{poly}}\log n$ (since we are going to take limit w.r.t. $n$, we view all other values independent of $n$ as constants; notice that $\ell$ is a fixed index here), such that $\left\{{\bm{v}}_{k}^{(\ell)}\right\}_{k=1}^{n}$ is a good sequence converging to $\left\{{\bm{z}}_{y}^{\prime(\ell)}\right\}_{y\in[c]}$ with parameter $\gamma^{(\ell)}$. Therefore, we have
@@
  • Next, we consider the layer-wise evolution. Corollary˜11 shows that for any $\ell\in[L]$,
@@
    Corollary˜11 also confirms that $\left\|{\bm{Z}}^{\prime(L)}{\bm{D}}^{1/2}{\bm{X}}\right\|>0$. Using an easy induction, we have
@@ Appendix C The Role of FFN: What Mappings are Great Mappings
The proof of Lemma˜12 is obvious.
@@
The proof of Lemma˜13 is obvious.
@@ Appendix D A Generalized Framework That is Independent of Data Generating
In the main paper, Theorem˜1 conditions on the specific data generating process used in Park et al. (2024). This is because we need to match the distribution of the attention connection to each token with their stationary distribution in the data. In this section, we show that, with a slightly stronger assumption on the attention map, a general result of the context-wise convergence can be derived.
@@
Notice that Definition˜6 is basically Definition˜3 but limits the function $f$ maps each node to only one node, instead of a set of nodes as in Definition˜3 (and in this case eq.˜5 automatically holds). Although this condition seems stronger, we note that following the same idea used in the main paper, that we can compose multiple attention maps into one, this still represents a large family of allowed attention maps.
Next, we prove a similar result as Theorem˜1 under Definition˜6 that is independent of input distribution.
@@
Using, Lemma˜7 which is also independent of DGP, we can prove the generalized results for a large family of attention maps by combining multiple attention maps that satisfies Definition˜6.